OpenStudy (anonymous):

There are values of t so that sin t=.3 and cos t=-.6. True or False. Is it true? Please explain.

OpenStudy (anonymous):

It's false :) And if you need help on that type of stuff, you also have google, my friend ;)

OpenStudy (anonymous):

sin^2t+cos^2=1. (0,3)^2+(-0,6)^2=1. 0,09+0,36=1 0,45=/1 False

OpenStudy (anonymous):

If I helped, a medal would be appreciated! :D